### v3

# Representational and Behavioral Stability of Truth in Large Language Models

Abstract
Large language models (LLMs) are increasingly used as information sources, yet small changes in semantic framing can destabilize their truth judgments. We propose P-StaT (Perturbation Stability of Truth), an evaluation framework for testing belief stability under controlled semantic perturbations in representational and behavioral settings via probing and zero-shot prompting. Across sixteen open-source LLMs and three domains, we compare perturbations involving epistemically familiar Neither statements drawn from well-known fictional contexts (Fictional) to those involving unfamiliar Neither statements not seen in training data (Synthetic). We find a consistent stability hierarchy: Synthetic content aligns closely with factual representations and induces the largest retractions of previously held beliefs, producing up to $32.7\%$ retractions in representational evaluations and up to $36.3\%$ in behavioral evaluations. By contrast, Fictional content is more representationally distinct and comparatively stable. Together, these results suggest that epistemic familiarity is a robust signal across instantiations of belief stability under semantic reframing, complementing accuracy-based factuality evaluation with a notion of epistemic robustness.
1 Introduction
Figure 1: Overview of the P-StaT framework. P-StaT evaluates the stability of an LLM $\mathcal{M}$’s beliefs by using a decision function $g$ to map statements $s$ to True or Not True under semantic assumptions $\Theta$. (a) In the baseline case $\Theta_{0}$, only True statements (purple) are labeled True, while False (green) and Neither (orange) statements are labeled Not True. Under a perturbed assumption set $\Theta$, a subset of Neither statements is treated as True. (b) The sets of True statements that $\mathcal{M}$ believes to be True in the baseline ($\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$) and perturbed ($\mathcal{B}^{\Theta}_{\mathrm{true}}$) cases are identified via $g$. (c) Comparing these belief sets reveals unstable beliefs that are either retracted ($\mathcal{R}$) or expanded ($\mathcal{E}$). Crucially, $g$ can be instantiated using either representation-based probing or behavioral zero-shot prompting, enabling direct connections between instability in latent space and behavioral instability.
Large language models (LLMs) are increasingly used as sources of information, but their behavior often blurs the line between knowledge and plausibility [2, 12]. While humans are expected to distinguish between True, False, and Neither-valued statements, it remains unclear whether LLMs form similarly structured internal representations, or whether such representations predict when truth judgments remain stable under changes in semantic framing [1, 26]. This gap is increasingly consequential as LLMs are deployed in high-stakes informational settings, where instability can undermine reliability and contribute to undesirable behaviors such as hallucinations [19, 16].
Existing work approaches this problem from two largely separate directions. Representation-based probing studies ask whether True and False statements occupy separable regions in activation space and whether “truth directions” transfer across LLMs or domains [20, 5, 23]. Behavioral studies document that small changes in prompting, context, or adversarial framing can substantially alter an LLM’s outputs, including its apparent confidence and self-consistency [29, 9, 18, 30]. While both lines of work reveal important failure modes, they are rarely connected through a shared experimental lens. Thus, we lack a unified framework for testing whether a candidate stability signal in latent space corresponds to stability in LLM behavior [13, 14].
We address this gap by proposing the Perturbation Stability of Truth framework, P-StaT, which links internal representations and LLM behavior through the lens of belief stability (Fig. 1). P-StaT asks how stable LLM truth judgments remain when semantic assumptions are systematically varied. This perspective is motivated by formal epistemology: under Leitgeb’s notion of $P$-stability, a belief system is epistemically well-formed only if established beliefs are preserved under small, principled changes in evidential context [17].
We operationalize semantic variation through perturbations $\Theta$ that modify which Neither statements are treated as compatible with truth. We distinguish epistemically familiar Neither statements (Fictional) from unfamiliar Neither statements (Synthetic) across three domains. Crucially, the same perturbation is instantiated both representationally, by retraining probes, and behaviorally, via zero-shot prompting.
We observe a consistent pattern: Unfamiliar Synthetic content induces the largest epistemic retractions in both representational and behavioral settings. By contrast, familiar Fictional content occupies more distinct regions of activation space and produces substantially fewer retractions. Together, these results identify epistemic familiarity as a cross-instantiation signal for when LLM truth judgments are most likely to destabilize under semantic reframing.
Contributions
  1. 1.
    We introduce a dataset of fictional statements across three domains (City Locations, Medical Indications, and Word Definitions) to enable controlled comparisons between epistemically familiar and unfamiliar Neither content.
  2. 2.
    We propose P-StaT, a perturbation-based stability framework grounded in epistemic notions of belief stability that supports both probing and zero-shot prompting.
  3. 3.
    Across sixteen open-source LLMs, we show that unfamiliar Synthetic statements induce the largest epistemic retractions both representationally and behaviorally, identifying epistemic familiarity as a key signal of stability.
2 Related Work
Our work lies at the intersection of three research threads: (i) representation-based probing of veracity, (ii) behavioral sensitivity to semantic context, and (iii) epistemic notions of belief stability.
Representational Probing of Veracity.
Probing examines which properties are recoverable from hidden representations of LLMs, providing insights into what LLMs encode beyond observable behavior [7, 15, 27]. Recent work has shown that True and False statements form separable clusters across LLMs and domains in activation space [20, 5]. The sAwMIL framework [23] extends this work by modeling True, False, and Neither statements as distinct representational directions. Related work on hallucination detection similarly suggests that hidden representations encode strong veracity signals even when outputs are wrong [12].
Behavioral Stability and In-Context Effects.
A separate line of research documents small changes in wording or framing can lead to large shifts in LLM behavior, including jailbreak vulnerabilities [30], sycophancy [25], and instability under paraphrasing [9]. Recent work has framed this brittleness in terms of behavioral consistency across interactions. For example, Li et al. introduce benchmarks for evaluating response stability in multi-turn settings [18]. Related work on in-context learning suggests that larger LLMs can override pretraining-induced semantics when presented with conflicting in-context labels, while smaller LLMs rely more heavily on prior knowledge [31]. Complementarily, Bigelow et al. propose a Bayesian account in which both prompting and activation steering modify latent beliefs [4].
Epistemic Belief Stability.
LLMs also exhibit systematic difficulties in tracking epistemic distinctions such as belief, knowledge, and factuality [26, 1]. Herrmann and Levinstein argue that the study of LLM beliefs lacks unified criteria for when internal representations should count as belief-like [14]. Formal epistemology offers a complementary perspective: Leitgeb’s notion of $P$-stability characterizes rational belief systems as those that preserve established beliefs under small, justified changes in evidential context [17].
By combining representational probing with behavioral perturbations through P-StaT, our work bridges prior studies of latent veracity structure and in-context sensitivity. We assess how beliefs respond to principled shifts in semantic assumptions, providing a unified view of epistemic stability across representations and behavior.
Dataset True False Synthetic Fictional Noise Examples
City
Locations
A: $1392$
N: $1376$
A: $1358$
N: $1374$
A: $876$
N: $876$
A: $350$
N: $350$
$795$
(T) The city of Surat is located in India.
(Fa) The city of Palembang is located in the Dominican Republic.
(S) The city of Norminsk is located in Jamoates.
(Fi) The city of Bikini Bottom is located in the Pacific Ocean.
Medical
Indications
A: $1439$
N: $1522$
A: $1523$
N: $1419$
A: $478$
N: $522$
A: $402$
N: $402$
$771$
(T) Pentobarbital is indicated for the treatment of insomnia.
(Fa) Vancomycin is not indicated for the treatment of lower respiratory tract infections.
(S) Alumil is indicated for the treatment of reticers.
(Fi) The Trump Virus is indicated for the treatment of Xenovirus Takis-B.
Word
Definitions
A: $1234$
N: $1235$
A: $1277$
N: $1254$
A: $1747$
N: $1753$
A: $1224$
N: $1224$
$1095$
(T) Hoagy is a synonym of an Italian sandwich.
(Fa) Decalogue is an astronomer.
(S) Dostab is a scencer.
(Fi) Snozzberry is a type of berry.
Table 1: Summary of datasets and statement types. Number of affirmative (A) and negated (N) statements across the three datasets, along with examples. Each dataset includes True, False, Synthetic, and Fictional statements, while Noise consists of randomly generated Gaussian activation vectors matched in dimensionality and distribution to the statement embeddings. Synthetic statements serve as Neither statements that were not seen during LLM training, i.e., $\mathcal{N}_{\mathrm{unf}}$, while Fictional statements are familiar Neither statements $\mathcal{N}_{\mathrm{fam}}$. A version of this table without the Fictional and Noise columns can be found in [23].
3 Methodology
We study how LLMs organize statements in activation space and whether this organization predicts the stability of beliefs under semantic reframing using P-StaT (Perturbation Stability of Truth), which applies the same perturbations in both probing and zero-shot evaluations.
3.1 Epistemic Familiarity and Neither Statements
Let $\mathcal{S}=\{s_{i}\}_{i=1}^{N}$ denote declarative statements drawn from factual domains with labels $y_{i}\in\{\texttt{True},\texttt{False},\texttt{Neither}\}$. Let $\mathcal{N}=\{s_{i}\in\mathcal{S}\mid y_{i}=\texttt{Neither}\}$ denote the set of Neither statements. While all statements in $\mathcal{N}$ lack real-world truth value, $\mathcal{N}$ is heterogeneous and can be partitioned into familiar (Fictional) and unfamiliar (Synthetic) subsets:
$\mathcal{N}=\mathcal{N}_{\mathrm{fam}}\cup\mathcal{N}_{\mathrm{unf}},\qquad\mathcal{N}_{\mathrm{fam}}\cap\mathcal{N}_{\mathrm{unf}}=\emptyset.$
Here, $\mathcal{N}_{\mathrm{fam}}$ contains well-known fictional entities likely present in training corpora, while $\mathcal{N}_{\mathrm{unf}}$ contains constructed entities intended to be absent from training data. For Fictional, we additionally distinguish canonically true and false subsets, $\mathcal{N}_{\mathrm{fam}}^{(T)}$ and $\mathcal{N}_{\mathrm{fam}}^{(F)}$.
We evaluate three domains that differ in how sharply truth and falsehood are delineated (City Locations, Medical Indications, Word Definitions; Table 1) [23, 20, 5]. True, False, and Synthetic statements originate in [23], while we introduce the new Fictional dataset. [^1] The new fictional dataset can be accessed at https://huggingface.co/datasets/samanthadies/representational_stability. Details on data construction and validation are in Supplementary Section B.
3.2 Veracity Representations
[…]
3.3 P-StaT: Perturbation Stability of Truth
P-StaT evaluates stability for a fixed LLM $\mathcal{M}$ by assuming the evaluation procedure
$g:(s_{i},\Theta)\mapsto\{\texttt{True},\texttt{Not True}\},$
where $\Theta$ encodes semantic assumptions about which statements should be treated as compatible with truth. The $g(\cdot,\Theta)$ interface is the core abstraction of P-StaT, enabling the same semantic perturbation to be instantiated in distinct evaluation settings.
Representational vs. Behavioral Stability.
In representational experiments, $g$ is implemented as $g_{\mathrm{p}}(s,\Theta)=h_{\Theta}(\phi_{\mathcal{M},l}(s))$, where $h_{\Theta}$ is a linear probe trained with labels induced by $\Theta$. In behavioral experiments, $g$ is implemented as $g_{\mathrm{zs}}(s,\Theta)$ via zero-shot prompting and a belief context $C_{\Theta}$. Both settings support the same perturbations, enabling direct comparison.
Baseline evaluation.
Under a baseline semantic interpretation $\Theta_{0}$, $g$ identifies the set of ground-truth True statements that are assigned True:
$\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}\;=\;\{\,s_{i}\mid y_{i}=\texttt{True},\;g(s_{i},\Theta_{0})=\texttt{True}\,\}.$
Perturbations.
A perturbation modifies the semantic assumptions encoded in $\Theta$. Let $\mathcal{N}_{\Theta}\subseteq\mathcal{N}$ denote the subset of Neither statements treated as compatible with truth under a perturbed interpretation. The perturbed interpretation $\Theta$ differs from $\Theta_{0}$ only in this reassignment and yields the perturbed belief set:
$\mathcal{B}^{\Theta}_{\mathrm{true}}\;=\;\{\,s_{i}\mid y_{i}=\texttt{True},\;g(s_{i},\Theta)=\texttt{True}\,\}.$
Thus, only the semantic interpretation of Neither content is altered.
Epistemic retractions.
Within P-StaT, we quantify stability by comparing the baseline and perturbed belief sets. Epistemic retractions, $\mathcal{R}\;=\;\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta}_{\mathrm{true}}$, capture ground-truth True statements that lose belief status under perturbation. While we also consider epistemic expansions, $\mathcal{E}\;=\;\mathcal{B}^{\Theta}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$, for completeness, epistemic retractions are a stronger signal of belief instability than expansions because they withdraw previously held beliefs [17].
Condition $\mathcal{N}_{\Theta}$ Probing: Training Labels Zero-shot: Belief Context $C_{\Theta}$
Baseline $\emptyset$ True vs. False+Synthetic+Fictional+Noise None
Synthetic $\mathcal{N}_{\mathrm{unf}}$ True+Synthetic vs. False+Fictional+Noise Synthetic
Fictional $\mathcal{N}_{\mathrm{fam}}$ True+Fictional vs. False+Synthetic+Noise Fictional
Fictional (T) $\mathcal{N}_{\mathrm{fam}}^{(T)}$ True+Fictional(T) vs. False+Synthetic+Fictional(F)+Noise Fictional (T)
Noise N/A True+Noise vs. False+Synthetic+Fictional N/A
Table 2: Perturbation conditions $\Theta$ and instantiations in P-StaT. Each condition corresponds to a semantic interpretation $\Theta$ defined by $\mathcal{N}_{\Theta}$, the Neither statements treated as compatible with truth. The same $\Theta$ is instantiated (i) representationally by retraining a probe with labels induced by $\mathcal{N}_{\Theta}$ and (ii) behaviorally by constructing a belief context $C_{\Theta}$ from the corresponding training statements. Noise is probing-only since it is a non-semantic control.
4 Experiments
We use P-StaT to apply the same perturbation $\Theta$ to both (i) probe-based evaluations over $\phi_{\mathcal{M},l}$ and (ii) zero-shot evaluations via belief context. We implement all experiments in Python using PyTorch [21], NumPy, SciPy, and scikit-learn [22], and rely on the HuggingFace Transformers ecosystem for LLM access and data handling [34]. All experiments were run on a university HPC cluster with NVIDIA H200 GPUs and required approximately $36$ GPU-hours (activation extraction and zero-shot evaluations), with probe training performed on CPU.
4.1 Data
We use the three domains and statement types defined in Section 3.1 (Table 1). Fictional is newly released with this work, and Noise is a set of randomly generated Gaussian vectors that match the dimensionality and distribution of the veracity representations and serve as a probing-only non-semantic control. Supplementary Section B contains additional information on data construction.
Approximately $55\%$ of the data is used for training, $20\%$ for calibration, and $25\%$ for testing (see Supplementary Tab. A2). The splits are shared across all experiments. We compute stability on the same held-out set of ground-truth True statements from the test split, so that differences in retractions are attributable only to changes in $\Theta$.
4.2 Activations
[…]
For descriptive analyses, we reduce each activation sequence to a single vector by selecting the final non-padding token. At the linguistic level, we compute rank–frequency curves over character bigrams aggregated across entity names for each statement type. At the representation level, we compute pairwise $1$-D Wasserstein distances between activation distributions, considering all activation dimensions. These measures reveal similarities between statement types at the linguistic level and in latent space.
4.3 Perturbation Conditions and Shared Protocol
For each perturbation $\Theta$ in Table 2, we instantiate $\Theta$ using only training data: probes are retrained with labels induced by $\Theta$, and belief contexts $C_{\Theta}$ are constructed from the corresponding training statements. We then evaluate on the same held-out set of ground-truth True test statements, computing $\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$ and $\mathcal{B}^{\Theta}_{\mathrm{true}}$ by applying $g(\cdot,\Theta_{0})$ and $g(\cdot,\Theta)$, and report epistemic retractions $\mathcal{R}=\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta}_{\mathrm{true}}$ (and expansions analogously).
4.3.1 Instantiation I: Probing over Activations
To implement $g(\cdot,\Theta)$ representationally, we train a linear probe that operates on the veracity representations $\phi_{\mathcal{M},l}(s)$. We use the sparse-aware multiple-instance learning probe (sAwMIL) [23], a multiclass probing method designed to extract reliable and transferable veracity directions from LLM activations. Unlike simpler probes such as the Mean Difference classifier [20], which assumes that truth and falsehood lie along a single axis, sAwMIL models True, False, and Neither as distinct directions and aggregates token-level representations using multiple-instance learning. As a max-margin method, sAwMIL yields stable decision boundaries, making differences across perturbations more reflective of genuine structure in the LLM’s geometry than noise from the probe. [^2] For comparison, results from the Mean Difference probe appear in Supplementary Section E.
For each condition in Table 2, we retrain the probe on $\mathcal{D}_{\mathrm{train}}$ with labels induced by $\Theta$, holding the architecture and hyperparameters fixed across conditions. Token-level representations are scaled using a standard scaler fit on the training set; bags are truncated to a fixed maximum size; and we perform a grid search over the regularization parameter $\mathcal{C}$ using three-fold cross-validation with mean average precision. [^3] Our code with all seeds and final hyperparameters is at https://github.com/samanthadies/representational_stability.
4.3.2 Instantiation II: Zero-shot Prompting with Belief Context
To implement $g(\cdot,\Theta)$ behaviorally, we use a zero-shot prompt. For each condition $\Theta$ (Tab. 2), we construct a belief context $C_{\Theta}$ from the corresponding training statements and insert it into each held-out ground-truth True test statement $s$.
To construct $C_{\Theta}$, we uniformly sample $K=100$ context statements from $\mathcal{N}_{\Theta}$ as specified in Table 2. We sample without replacement.
For each ground-truth True test statement, we use the following prompt:
[optional belief context $C_{\Theta}$]
Is the following statement correct?
 [statement $s$]
a. The statement is true.
b. The statement is false.
c. The statement is neither true nor false.
The final answer is
For chat-tuned LLMs, we place the same content into the LLM’s chat template and define the model’s response as the highest-probability next token among the answer labels $\{a,b,c\}$ at the final prompt position.
We compute next-token probabilities for $a$, $b$, and $c$ at the final position and predict
$g_{\mathrm{zs}}(s,\Theta)=\mathop{\mathrm{arg\,max}}\limits_{\ell\in\{a,b,c\}}p(\ell\mid s,\Theta),$
with temperature set to zero. [^4] We also track the probability mass corresponding to all non-label tokens, but define $g$ exclusively over $\{a,b,c\}$ rather than parsing free-form textual outputs. We treat option $a$ as True and options $\{b,c\}$ as Not True. All zero-shot evaluations are therefore deterministic, and reported results aggregate outcomes across LLMs rather than multiple random seeds.
5 Results
Figure 2: Linguistic vs. representational structure of Neither statements. For (a),(b) City Locations, (c),(d) Medical Indications, and (e),(f) Word Definitions, the left column shows normalized character bigram rank–frequency curves for True (green), False (red), Synthetic (yellow), and Fictional (blue) statements, and the right column shows pairwise Wasserstein distances between activation distributions. Fictional content exhibits a domain-dependent decoupling between linguistic form and latent-space organization, indicating that veracity representations reflect both epistemic context and linguistic form.
Table 3: Epistemic expansions $\mathcal{E}$ and retractions $\mathcal{R}$ under probing label perturbations. Counts (percentages) of beliefs that remain stable or undergo expansions $\mathcal{E}$ or retractions $\mathcal{R}$ under Synthetic (yellow), Fictional (gray), Fictional(T) (blue), and Noise (red) perturbations. Synthetic perturbations consistently produce the highest rate of epistemic retractions, indicating that epistemically unfamiliar content induces the most instability.
Table 4: Epistemic expansions $\mathcal{E}$ and retractions $\mathcal{R}$ under perturbed zero-shot experiments. Counts (percentages) of beliefs that remain stable or undergo expansions $\mathcal{E}$ or retractions $\mathcal{R}$ under Synthetic (yellow), Fictional (gray), and Fictional(T) (blue) perturbations. Despite domain variation, Synthetic perturbations induce the highest retraction rates in all cases, mirroring the hierarchy observed under probing in Table 3.
We report (i) the descriptive structure of Neither statements, (ii) representational stability under probing, and (iii) behavioral stability under zero-shot prompting. LLM-level results are in Supplementary Section D.
5.1 Linguistic and Representational Structure of Neither Statements
We begin by characterizing how familiar and unfamiliar Neither statements differ. At the linguistic level, True, False, and Synthetic statements exhibit nearly identical normalized bigram rank–frequency curves (Figure 2(a),(c),(e)). Fictional statements, by contrast, show a slower decay reflecting stylistic patterns characteristic of narrative text. The differences in Fictional bigram distributions are most visible in Word Definitions and least in City Locations.
We next test whether these linguistic differences explain latent-space structure by comparing activation distributions across statement types using pairwise Wasserstein distances (Figure 2(b),(d),(f); LLM-level heatmaps appear in Supplementary Section D.1). Across domains, True and False occupy nearby regions of activation space, and Synthetic remains relatively close to factual content despite its lack of real-world referents. By contrast, Fictional statements are consistently more representationally separated from factual content, forming a distinct cluster that is not well-explained by bigram statistics alone.
Taken together, these findings indicate a decoupling between linguistic and latent-space similarity. Although linguistic divergence of Fictional content varies by domain, representational separation remains consistently pronounced. Further, the domain with the largest Fictional separation linguistically (Word Definitions) is the domain with the smallest representational distance. The geometry of LLM activations thus reflects both linguistic form and context.
5.2 Representational Stability under Probing Perturbations
We next evaluate P-StaT representational stability by retraining sAwMIL under perturbed interpretations $\Theta$ and measuring epistemic retractions on held-out ground-truth True statements. Table 3 summarizes prediction changes across perturbation conditions aggregated over all LLMs. LLM-level results and results for the Mean Difference probe appear in Supplementary Sections D.2 and E.
Across all three domains, Synthetic perturbations induce the highest rate of epistemic retractions: $2.1\%$ in City Locations, $4.5\%$ in Medical Indications, and $32.7\%$ in Word Definitions. These rates are substantially larger than those produced by Fictional and Fictional(T) perturbations, establishing a clear perturbation hierarchy in which epistemically unfamiliar Neither content most strongly destabilizes learned veracity boundaries. Synthetic perturbations also induce the largest epistemic expansions across domains.
Retraction rates also vary systematically by domain. City Locations is highly stable, Medical Indications shows moderate instability, and Word Definitions is markedly more fragile. These results again indicate that representational stability is governed by epistemic familiarity, given that City Locations are likely richly represented in training corpora, while Medical Indications are common but specialized, and Word definitions are the most semantically flexible.
5.3 Behavioral Stability under Zero-Shot Perturbations
We next evaluate P-StaT behavioral stability by applying the same perturbations $\Theta$ via belief context in zero-shot prompting and measuring epistemic retractions. Table 4 reports belief changes across LLMs, with LLM-level results in Supplementary Section D.3. Overall belief changes are larger than in probing experiments, reflecting the greater flexibility of behavioral responses compared to linear reclassification over fixed representations.
Despite this increase, the perturbation hierarchy observed representationally also appears behaviorally. The Synthetic perturbation induces the highest epistemic retraction rate in every domain, with $36.3\%$ in City Locations, $27.1\%$ in Medical Indications, and $21.4\%$ in Word Definitions. Fictional and Fictional(T) also induce substantial retractions, but consistently fewer than Synthetic within each dataset.
However, domain ordering differs from the probing setting. City Locations has the highest zero-shot retraction rates, while Word Definitions has the lowest. This reversal likely reflects interactions between belief stability and behavioral accuracy. Specifically, zero-shot retractions conflate belief changes with the number of baseline beliefs available to retract (i.e., a statement must be considered True in the baseline case in order to become a retraction). In contrast, probing isolates changes in the decision boundary under fixed representations.
Our results show that epistemic familiarity governs stability not only in activation space, but also in LLM behavior. Unfamiliar Synthetic content systematically destabilizes LLM beliefs under semantic perturbation, whereas familiar Fictional content produces smaller and more context-dependent effects.
6 Discussion
We provide a unified perspective on how LLMs organize and maintain truth judgments under semantic perturbation. Across LLMs, domains, and evaluation settings, Neither statements play a central role in belief stability: unfamiliar Synthetic statements consistently induce the largest epistemic retractions, while familiar Fictional statements are more stable. Our findings indicate that belief stability in LLMs is governed as much by epistemic familiarity as by linguistic form.
At the representational level, we observe a decoupling between linguistic similarity and latent-space organization. Fictional content exhibits domain-dependent linguistic variation that does not align with distance in representation space. This mismatch suggests that veracity representations encode higher-level epistemic context learned during training rather than lexical statistics alone. Through the lens of P-StaT, these representational differences translate into systematic differences in stability under semantic perturbation. Expanding the semantic definition of truth to include unfamiliar Synthetic content reliably destabilizes previously held beliefs, while perturbations involving familiar Fictional content are better tolerated. These results hold in both representational and behavioral instantiations, suggesting that latent-space organization captures meaningful epistemic structure.
More broadly, P-StaT complements accuracy-based factuality benchmarks by shifting focus from correctness under fixed prompts to stability under principled semantic shifts. By emphasizing epistemic retractions, it targets a particularly consequential form of instability. This perspective aligns with formal accounts of rational belief change, such as Leitgeb’s notion of $P$-stability, which requires that established beliefs be preserved under justified changes in evidential context [17]. From this viewpoint, the sensitivity of LLMs to Synthetic perturbations highlights a structural limitation: distributional plausibility alone does not ensure a stable internal organization of truth, falsity, and indeterminacy. Together, these findings suggest that evaluating belief stability under controlled semantic perturbations offers a principled way to probe the epistemic organization of LLMs beyond what can be inferred from accuracy alone.
7 Conclusion
We show that epistemic familiarity is a key determinant of belief stability under semantic reframing. By linking internal representations and behavioral responses under matched perturbations, P-StaT reveals that unfamiliar Synthetic content consistently induces the largest epistemic retractions across LLMs and domains. These results demonstrate that evaluating belief stability under controlled perturbations of Neither content provides a principled complement to accuracy-based factuality metrics and a step toward more epistemically reliable language models.
8 Limitations
Our perturbations focus on a specific notion of epistemic familiarity and a limited set of factual domains. While this design enables controlled comparisons across representational and behavioral settings, it does not exhaustively cover all types of semantic variation. Extending P-StaT to other forms of epistemic ambiguity, such as disputed claims, probabilistic beliefs, or evolving facts, would further test its generality. In addition, our analysis considers fixed LLM parameters. Although P-StaT isolates how semantic perturbations interact with existing internal representations, it does not address how belief stability may change in settings where representations themselves evolve over time. Finally, retraction rates depend on accuracy on True statements, so our notion of stability reflects both robustness to perturbation and an LLM’s propensity to assert truth in the baseline case. While we emphasize epistemic retractions as a primary signal of instability, other applications may require alternative notions of stability.
9 Ethical Considerations
Our study examines the conditions under which perturbations can systematically destabilize what an LLM considers true. For example, we find that unfamiliar synthetic content induces the largest epistemic retractions, producing up to 32.732.7% retractions in representational evaluations and up to 36.336.3% in behavioral evaluations. This could inform efforts to undermine LLM reliability and has implications for trust in LLMs.
While our framework could be misused to deliberately destabilize LLM beliefs, our intent is diagnostic rather than adversarial: P-StaT is designed to identify epistemic vulnerabilities in order to inform more robust evaluation and LLM design. We do not propose interventions or belief manipulation techniques, and all experiments are conducted on fixed, open-source LLMs in offline settings. More broadly, our findings highlight a structural limitation of current LLMs rather than a prescription for exploiting it. P-StaT is intended solely for research and diagnostic evaluation of pretrained LLMs and is not designed for deployment, belief steering, or real-world decision-making systems.
9.1 Data availability
We use the True, False, and Synthetic statements available at https://huggingface.co/datasets/carlomarxx/trilemma-of-truth under a CC-BY-4.0 license. Fictional statements are available at https://huggingface.co/datasets/samanthadies/representational_stability. The code used to generate the Noise activations can be found at https://github.com/samanthadies/representational_stability.
9.2 Code availability
We release the code used to generate activations, generate the Noise activations, and run the P-StaT stability experiments under an MIT License at https://github.com/samanthadies/representational_stability. ChatGPT and Copilot were used to help write initial versions of experiment and plotting scripts, and helped clean and comment the code. All AI-generated content was reviewed and verified by the authors.
All models used in this work are publicly available for research use under their respective licenses (MIT License for sAwMIL; Gemma for Gemma-7b, Gemma-7b-it, Gemma-2-9b, and Gemma-2-9b-it; llama3.1 for Llama-3.1-8b and Llama-3.1-8b-Instruct; llama3.2 for Llama-3.2-3b, Llama-3.2-3b-Instruct; llama3 for Llama3-Med42-8b; Bio-Medical-Llama-3-8b LLM License for Bio-Medical-Llama-3-8b; apache-2.0 for Mistral-7B-v0.3, Mistral-7B-Instruct-v0.3, Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-14B, and Qwen2.5-14B-Instruct). Marks and Tegmark did not specify a license for the Mean Difference probe [20].
Acknowledgments
We thank Hannes Leitgeb and Branden Fitelson for discussions on $P$-stability and how it might be related to epistemic uncertainty in LLMs. We also thank Zohair Shafi and Moritz Laber for their feedback and discussions on methodological and empirical portions of this work.
Funding
This material was sponsored by the Government of the United States under Contract Number FA8702-15-D-0002. The view, opinions, and/or filings contained in this material are those of the author(s) and should not be construed as an official position, policy, or decision of the Government of the United States or Carnegie Mellon University or the Software Engineering Institute unless designated by other documentation.
Competing interests
The authors declare no competing interests.
Appendix A Notation
We summarize the mathematical notation used throughout the manuscript in Table A1.
Symbol Description
$\mathcal{M}$ A fixed large language model (LLM).
$l$ Layer index used for activation extraction.
$\mathcal{S}=\{s_{i}\}_{i=1}^{N}$ Set of $N$ natural-language statements.
$y_{i}$ Ground-truth veracity label of $s_{i}$, $y_{i}\in\{\texttt{True},\texttt{False},\texttt{Neither}\}$.
$\mathcal{N}$ Set of all Neither statements: $\mathcal{N}=\{s_{i}\mid y_{i}=\texttt{Neither}\}$.
$\mathcal{N}_{\mathrm{fam}},\ \mathcal{N}_{\mathrm{unf}}$ Partition of $\mathcal{N}$ into epistemically familiar (Fictional) and unfamiliar (Synthetic) subsets.
$\mathcal{N}_{\mathrm{fam}}^{(T)},\ \mathcal{N}_{\mathrm{fam}}^{(F)}$ Canonically true and false subsets of familiar fictional statements (used in Fictional(T)).
$\phi_{\mathcal{M},l}$ Representation map from a statement to its layer-$l$ hidden representations under $\mathcal{M}$.
$\mathbf{z}_{i}^{(l)}$ Layer-$l$ representation of statement $s_{i}$, $\mathbf{z}_{i}^{(l)}=\phi_{\mathcal{M},l}(s_{i})$.
$\mathcal{D}$ Dataset of statements, representations, and labels: $\mathcal{D}=\{(s_{i},\mathbf{z}_{i}^{(l)},y_{i})\}_{i=1}^{N}$.
$\mathcal{D}_{\mathrm{train}},\ \mathcal{D}_{\mathrm{test}}$ Training and test splits of $\mathcal{D}$.
$\Theta$ Semantic interpretation specifying which Neither statements are treated as compatible with truth.
$\mathcal{N}_{\Theta}\subseteq\mathcal{N}$ Subset of Neither statements treated as compatible with truth under $\Theta$.
$g(\cdot,\Theta)$ Evaluation function mapping a statement and interpretation to a binary judgment: $g(s,\Theta)\in\{\texttt{True},\texttt{Not True}\}$.
$g_{\mathrm{p}}$ Representational instantiation of $g$ via a probe over $\phi_{\mathcal{M},l}(s)$.
$g_{\mathrm{zs}}$ Behavioral instantiation of $g$ via zero-shot prompting with belief context $C_{\Theta}$.
$C_{\Theta}$ The belief context inserted into a zero-shot prompt to instantiate perturbation $\Theta$.
$\Theta_{0}$ Baseline semantic interpretation (no Neither treated as true).
$\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$ Baseline belief set: $\{s_{i}\mid y_{i}=\texttt{True},\ g(s_{i},\Theta_{0})=\texttt{True}\}$.
$\mathcal{B}^{\Theta}_{\mathrm{true}}$ Belief set under perturbed interpretation $\Theta$.
$\mathcal{R}$ Epistemic retractions: $\mathcal{R}=\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta}_{\mathrm{true}}$.
$\mathcal{E}$ Epistemic expansions: $\mathcal{E}=\mathcal{B}^{\Theta}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$.
Table A1: Notation. Summary of symbols used throughout the manuscript.
Appendix B Data
B.1 Data Generation
We use statements from the City Locations, Medical Indications, and Word Definitions datasets introduced in [23]. City statements take the form “The city of [city] is (not) located in [country],” (omitting “The city of” when redundant). Medical statements follow “[drug] is (not) indicated for the treatment of [disease/condition].” Word Definition statements draw from three templates: “[word] is (not) a [instanceOf],” “[word] is (not) a type of [typeOf],” and “[word] is (not) a synonym of [synonym].” No personal data, identifying information, or user-generated content is included.
B.1.1 True, False, and Synthetic Statements
We take the True, False, and Synthetic statements from the datasets introduced in [23]. All statements are constructed with both affirmative and negated forms. Synthetic entities are generated using a Markov-chain–based name generator (namemaker [^5] https://github.com/Rickmsd/namemaker.) and undergo multi-stage filtering, including database checks, model tagging, and web-search validation, to ensure no accidental overlap with real entities. Validated names are then paired to form grammatically coherent but semantically meaningless statements that follow each template. Because Synthetic entities do not exist and cannot have appeared in training corpora, LLMs have no basis for assigning them a truth value. Accordingly, these statements function as Neither cases: unknown claims for which belief should be suspended rather than confidently classified as true or false.
B.1.2 Fictional Statements
In addition to Synthetic statements, which represent unseen and unknown claims, we construct new sets of Fictional statements for all three domains. Fictional statements also function as Neither statements in our experiments as they reference entities that do not exist in the real world and therefore lack real-world truth value. However, unlike Synthetic statements, many Fictional entities are likely to have appeared in LLM training corpora. [^6] For later analyses, we additionally annotate fictional statements with their within-universe factual status (Fictional (T) or Fictional (F)), but this labeling is not used in the primary True vs. Not True classification tasks. As such, they represent a complementary form of Neither: claims that an LLM may recognize, but that still lie outside the true–false axis relevant to factual grounding.
To ensure that Fictional statements remain genuinely non-factual, all terms were validated to exclude any real-world overlap, and fictional lexical items appearing in any natural language were excluded to prevent misinterpretation by multilingual LLMs. Fictional statements were then constructed using the same templates as the True, False, and Synthetic statements, including both affirmative and negated forms.
Fictional City Locations.
[…]
Fictional Medical Indications.
Fictional drug and disease statements are drawn from (1) NeoEncyclopedia Wiki [10, 11]; (2) ChemEurope’s List of Fictional Medicines and Drugs [6]; and (3) The Thackery T. Lambshead Pocket Guide to Eccentric & Discredited Diseases [28]. Drug–disease pairs are included when a treatment relationship exists according to the fictional source.
Fictional Word Definitions.
Fictional lexical items are compiled from (1) Gobblefunk [3]; (2) Dothraki [24]; and (3) Na’vi [8]. Dothraki and Na’vi have formal linguistic structure, whereas Gobblefunk is a playful neologistic extension of English.
B.1.3 Noise
The Noise statements contain no linguistic content. We generate $n_{\mathrm{noise}}=0.10\cdot|\mathcal{D}|$ random activation sequences by sampling from a multivariate Gaussian with per-feature mean, standard deviation, and sequence-length distribution matched to the LLM activations. These distributionally consistent but non-semantic sequences serve as a control, allowing us to test whether observed representational differences arise from semantic content or from statistical variation in activation space.
B.2 Data Splits for Probing Experiments
Dataset Train Calibration Test Total
City Locations $4746$ $(0.54)$ $1772$ $(0.20)$ $2229$ $(0.25)$ $8747$ $(1.00)$
Medical Indications $4636$ $(0.55)$ $1721$ $(0.20)$ $2121$ $(0.25)$ $8478$ $(1.00)$
Word Definitions $6488$ $(0.54)$ $2514$ $(0.21)$ $3041$ $(0.25)$ $12043$ $(1.00)$
Table A2: Dataset splits. The number of statements used in training, calibration, and testing of the probe. The proportion of total statements is reported in parentheses.
Table A2 summarizes the partitions used for all experiments. Each dataset is split exclusively into training, calibration, and test sets to prevent data leakage. Approximately $55\%$ of statements are used for training, $20\%$ for calibration, and $25\%$ for testing. We use identical splits in all conditions.
Appendix C LLMs
Table A3 lists the sixteen open-source LLMs used in our experiments. The set spans four major model families, Gemma, Llama, Mistral, and Qwen, with approximately $3$ billion to $15$ billion parameters and release dates between February and September $2024$. For each family, we include both base (pre-trained) and chat-tuned variants. Together, these LLMs provide a representative cross-section of current decoder-only architectures varying in scale, origin, and training objectives.
Official Name Short Name Type # Decoders # Parameters Best Layer Release Date Source
Gemma-77b gemma-7b Base $28$ $8.54$ B C: $14$, M: $19$, W: $17$ Feb $21$, $2024$ Google
Gemma-22-99b gemma-2-9b Base $26$ $9.24$ B C: $24$, M: $25$, W: $23$ Jun $27$, $2024$ Google
Llama-33-88b llama-3.1-8b Base $32$ $8.03$ B C: $18$, M: $17$, W: $17$ Jul $23$, $2024$ Meta
Llama-3.23.2-33b llama-3.2-3b Base $28$ $3.21$ B C: $16$, M: $17$, W: $15$ Sep $25$, $2024$ Meta
Mistral-77B-v0.30.3 mistral-7B-v0.3 Base $32$ $7.25$ B C: $18$, M: $17$, W: $18$ May $22$, $2024$ Mistral AI
Qwen2.52.5-77B qwen-2.5-7b Base $28$ $7.62$ B C: $18$, M: $19$, W: $17$ Sep $19$, $2024$ Alibaba Cloud
Qwen2.52.5-1414B qwen-2.5-14b Base $38$ $14.80$ B C: $30$, M: $31$, W: $30$ Sep $19$, $2024$ Alibaba Cloud
Gemma-77b-it _gemma-7b Chat $28$ $8.54$ B C: $19$, M: $19$, W: $17$ Feb $21$, $2024$ Google
Gemma-22-99b-it _gemma-2-9b Chat $26$ $9.24$ B C: $27$, M: $26$, W: $25$ Jul $27$, $2024$ Google
Llama-3.23.2-33b-Instruct _llama-3.2-3b Chat $28$ $3.21$ B C: $16$, M: $19$, W: $18$ Sep $25$, $2024$ Meta
Llama-3.13.1-88b-Instruct _llama-3.1-8b Chat $32$ $8.03$ B C: $18$, M: $19$, W: $18$ Jul $23$, $2024$ Meta
Llama33-Med4242-88b _llama-3-8b-med Chat $32$ $8.03$ B C: $18$, M: $16$, W: $15$ Aug $12$, $2024$ M42 Health
Bio-Medical-Llama-33-88b _llama-3-8b-bio Chat $32$ $8.03$ B C: $18$, M: $19$, W: $18$ Aug $11$, $2024$ Contact Doctor
Mistral-77b-Instruct-v0.30.3 _mistral-7B-v0.3 Chat $32$ $7.25$ B C: $19$, M: $21$, W: $18$ May $22$, $2024$ Mistral AI
Qwen2.52.5-77B-Instruct _qwen-2.5-7b Chat $28$ $7.62$ B C: $19$, M: $21$, W: $18$ Aug $18$, $2024$ Alibaba Cloud
Qwen2.52.5-1414B-Instruct _qwen-2.5-14b Chat $38$ $14.80$ B C: $31$, M: $34$, W: $30$ Aug $18$, $2024$ Alibaba Cloud
Table A3: LLMs used in the stability experiments. We list the official names of the LLMs according to the HuggingFace repository [34]. We further specify the shortened name we use to refer to each of the LLMs, whether it is the base, pre-trained LLM or a chat-tuned version, the number of decoders, the number of parameters, the release date, and the source of the LLM. Finally, we report the layers with the best separation between True and Not True statements for the City Locations (“C”), Medical Indications (“M”), and Word Definitions (“W”) datasets. The LLMs are publicly available through HuggingFace [34].
Appendix D LLM-level Results
The main text reports LLM-agnostic stability patterns aggregated across sixteen LLMs. Here we provide the corresponding LLM-level results.
D.1 Linguistic and Representational Structure of Neither Statements (by LLM)
Figures A1-A16 show the pairwise activation distance matrices for all sixteen LLMs. Three general representational patterns emerge. The first, observed in _gemma-2-9b (Fig. A1) and gemma-2-9b (Fig. A10), shows Fictional and Synthetic statements clustering near True and False statements, with Noise forming a distinct outlier. The second, present in _gemma-7b (Fig. A2), gemma-7b (Fig. A11), _qwen-2.5-14b (Fig. A8), qwen-2.5-14b (Fig. A15), and _qwen-2.5-7b (Fig. A9), exhibits Synthetic statements close to True and False, Fictional statements clearly separated, and Noise positioned slightly closer to the True/False/Synthetic cluster. The third, seen in the remaining nine LLMs, features Synthetic statements aligned with True and False, while both Fictional and Noise statements occupy distinct and distant regions. Except for _qwen-2.5-7b (which follows the second pattern; Fig. A9) and qwen-2.5-7b (the third; Fig. A16), base and chat versions are qualitatively similar.
Figure A1: Wasserstein distance between activations for _gemma-2-9b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Noise has distinct representations, but Fictional and Synthetic statements are represented similarly to True and False statements and each other.
Figure A2: Wasserstein distance between activations for _gemma-7b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements are represented distinctly from all other statements.
Figure A3: Wasserstein distance between activations for _llama-3-8b-med. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
Figure A4: Wasserstein distance between activations for _llama-3-8b-bio. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
Figure A5: Wasserstein distance between activations for _llama-3.1-8b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
Figure A6: Wasserstein distance between activations for _llama-3.2-3b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
Figure A7: Wasserstein distance between activations for _mistral-7B-v0.3. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
Figure A8: Wasserstein distance between activations for _qwen-2.5-14b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements are represented distinctly from all other statements.
Figure A9: Wasserstein distance between activations for _qwen-2.5-7b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements are represented distinctly from all other statements.
Figure A10: Wasserstein distance between activations for gemma-2-9b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Noise has distinct representations, but Fictional and Synthetic statements are represented similarly to True and False statements and each other.
Figure A11: Wasserstein distance between activations for gemma-7b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements are represented distinctly from all other statements.
Figure A12: Wasserstein distance between activations for llama-3-8b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
Figure A13: Wasserstein distance between activations for llama-3.2-3b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
Figure A14: Wasserstein distance between activations for mistral-7B-v0.3. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
Figure A15: Wasserstein distance between activations for qwen-2.5-14b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions. Synthetic statements are represented similarly to True and False statements, while Fictional statements are represented distinctly from all other statements.
Figure A16: Wasserstein distance between activations for qwen-2.5-7b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements and Noise are represented distinctly from all other statements.
D.2 Representational Stability under Probing Perturbations (by LLM)
Figure A17 shows how the sAwMIL decision boundaries change under each perturbation, measured by cosine similarity to the baseline True vs. Not True direction and by the associated bias shift. Consistent with the aggregate flip rates (Table 3), Synthetic produces the largest boundary changes across domains, while Fictional, Fictional(T), and Noise yield smaller deviations.
Different LLM families exhibit different degrees of susceptibility to these perturbations. Chat-tuned variants (denoted by leading underscores) tend to exhibit somewhat larger rotations and bias shifts than their base models. An exception is gemma-7b, which shows unusually large shifts in the Word Definitions domain.
Figures A18A20 break down epistemic retractions $\mathcal{R}$ and expansions $\mathcal{E}$ by LLM. Chat-tuned LLMs tend to produce more expansions, while Base models tend to exhibit more retractions. However, these tendencies do not hold uniformly, and overall differences across LLM families are smaller than differences across perturbation types.
Figure A17: Changes in the probe decision boundary under perturbations. Cosine similarity (left column) and bias difference (right column) between the baseline True vs. Not True probe and probes retrained under label perturbations for the (a,b) City Locations, (c,d) Medical Indications, and (e,f) Word Definitions datasets. Each heatmap shows results for sixteen LLMs (columns) and five perturbation conditions (rows). LLMs with leading underscores are Chat models, while those without are Base models. Higher cosine similarity indicates smaller rotations of the learned decision boundary, while bias difference reflects shifts in intercept. Across datasets, probes retrained with the Synthetic perturbation show the largest deviation from the original, particularly in cosine similarity.
Figure A18: Stability of probe predictions under belief context perturbations for City Locations data. Bar plots show for each LLM (x-axis) how often the probe induces epistemic expansions and retractions when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Base models exhibit more retractions than the Chat models, and the Chat models induce more expansions.
Figure A19: Stability of probe predictions under belief context perturbations for Medical Indications data. Bar plots show for each LLM (x-axis) how often the probe induces epistemic expansions and retractions when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Synthetic perturbation leads to the most instability, and the Fictional and Fictional (T) perturbations result in almost no flips.
Figure A20: Stability of probe predictions under belief context perturbations for Word Definitions data. Bar plots show for each LLM (x-axis) how often the probe induces epistemic expansions and retractions when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Synthetic perturbation leads to the most instability, with some LLMs retracting over $50\%$ of their originally True statements.
D.3 Behavioral Stability under Zero-Shot Perturbations (by LLM)
Figures A21A23 break down epistemic retractions $\mathcal{R}$ and expansions $\mathcal{E}$ by LLM. Opposite to the behavior seen in the representational instantiation, Base LLMs tend to produce more expansions, while Chat-tuned variants tend to exhibit more retractions.
Figure A21: Stability of zero-shot beliefs under belief context perturbations for City Locations data. Bar plots show for each LLM (x-axis) how often the zero-shot instantiation induced epistemic expansions and retractions when retrained under three perturbations: (a) Synthetic, (b) Fictional, and (c) Fictional(T). Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Chat models exhibit more retractions than the Base models.
Figure A22: Stability of zero-shot beliefs under belief context perturbations for Medical Indications data. Bar plots show for each LLM (x-axis), how often the zero-shot instantiation induced epistemic expansions and retractions when retrained under three perturbations: (a) Synthetic, (b) Fictional, and (c) Fictional(T). Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Chat models exhibit more retractions, while the Base models exhibit more expansions.
Figure A23: Stability of zero-shot beliefs under belief context perturbations for Word Definitions data. Bar plots show for each LLM (x-axis), how often the zero-shot instantiation induced epistemic expansions and retractions when retrained under three perturbations: (a) Synthetic, (b) Fictional, and (c) Fictional(T). Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Chat models exhibit more retractions, while the Base models exhibit more expansions.
Appendix E Exploring the Mean Difference Probe
We repeated the representation-based perturbation experiments using the Mean Difference probe proposed by Marks and Tegmark [20] to supplement the sAwMIL results. Mean Difference estimates a “truth direction” by taking the vector difference between the mean activation of True statements and that of False statements, optionally scaled by the inverse covariance matrix of the data. This approach is inherently sensitive to differences in the centroids and covariance structure of the data, which leads to strong instability in the learned decision boundary when Neither statements are included alongside True and False examples. The Mean Difference probe shows considerably greater variability across LLMs than sAwMIL (Fig. A25). While sAwMIL yields consistent decision boundary rotation corresponding to specific perturbations (see Fig. A17, particularly the Synthetic perturbation), the Mean Difference probe exhibits near-orthogonal boundary shifts for certain LLMs regardless of perturbation. In addition, Table A24 shows that, unlike with sAwMIL, the Fictional perturbation produces the largest epistemic retractions across all three datasets, and the Word Definitions dataset exhibits the fewest total retractions. We note, however, that the Synthetic perturbation still produces the most epistemic expansions across domains, consistent with the sAwMIL results. We interpret these discrepancies as artifacts of the Mean Difference probe’s reliance on dataset centroids: when statement activations are well separated, as with Fictional statements, class-label perturbations can induce disproportionately large changes in the estimated decision boundary. This instability reflects probe sensitivity rather than genuine representational instability in the LLMs. Accordingly, the Mean Difference probe is less well suited for quantifying stability within P-StaT than sAwMIL.
Figure A24: Epistemic expansions $\mathcal{E}$ and retractions $\mathcal{R}$ under probing label perturbations for the Mean Difference Probe. Counts (and percentages) of beliefs that remain stable or lead to expansions $\mathcal{E}$ and retractions $\mathcal{R}$ across Synthetic (yellow), Fictional (gray), Fictional(T) (blue), and Noise (red) perturbations for each dataset. While the Fictional perturbation induces the most epistemic retractions $\mathcal{R}$, the Synthetic perturbation induces the most epistemic expansions $\mathcal{E}$.
Figure A25: Change in Mean Difference decision boundaries under perturbations. Cosine similarity (left column) and bias difference (right column) between the baseline True vs. Not True probe and probes retrained under label perturbations for the (a,b) City Locations, (c,d) Medical Indications, and (e,f) Word Definitions datasets. Each heatmap shows results for sixteen LLMs (columns) and five perturbation conditions (rows). LLMs with leading underscores are Chat models, while those without are Base models. Higher cosine similarity indicates smaller rotations of the learned decision boundary, while bias difference reflects shifts in intercept. Certain LLMs lead to near orthogonal perturbed decision boundaries across all perturbation types, suggesting that, unlike sAwMIL, the probe is highly sensitive to differences in the distributions of the underlying activations.